# EFRame: Deeper Reasoning via Exploration-Filter-Replay Reinforcement Learning Framework

### v2

@@ Abstract
Recent advances in reinforcement learning (RL) have significantly enhanced the reasoning capabilities of large language models (LLMs). Group Relative Policy Optimization (GRPO), an efficient variant of PPO that lowers RL’s computational cost, still faces limited exploration, low sample efficiency and instability, constraining its performance on complex reasoning tasks. To address these limitations, we introduce EFRame, an Exploration-Filter-Replay framework that systematically augments GRPO along three critical dimensions. EFRame performs additional rollouts to explore high-quality trajectories, applies online filtering to eliminate low-quality samples that introduce noise and variance, and leverages experience replay to repeatedly exploit rare but informative samples. EFRame establishes a complete and stable learning cycle, guiding the model through a structured transition from exploration to convergence. Our experiments across a variety of reasoning benchmarks demonstrate that EFRame not only improves the robustness and efficiency of training, but also enables access to deeper reasoning capabilities that remain unattainable under vanilla GRPO. Furthermore, EFRame enables a more fine-grained categorization of training samples, allowing for a deeper analysis of how different types of samples contribute to the learning process in RL. Our code is avail able at https://github.com/597358816/EFRame.
@@ 1 Introduction
Previous studies have shown that both positive and negative samples contribute to RLVR training Zhu et al. (2025). With the support of the EFRame, we categorize training samples to understand their roles in learning better:
  • Regular samples refer to both correct and incorrect responses from prompts of moderate difficulty. They form the foundation of the reward signal, helping to improve model performance while ensuring training stability.
  • High-quality samples consist of correct responses obtained through the exploration of difficult prompts. They are crucial for activating deeper model knowledge and enhancing the performance ceiling, but their effective use requires experience replay to avoid instability caused by their low sampling probability.
  • Low-quality samples refer to the numerous incorrect responses generated from challenging prompts. These samples offer minimal learning value and may even harm the training process by accelerating convergence to suboptimal solutions.
EFRame provides a systematic solution to the inherent trade-offs among training stability, efficiency, and exploration capacity in reinforcement learning. By enabling the model to effectively identify and leverage deeper latent knowledge, EFRame enhances the expressiveness and learning potential of large language models. Through the coordinated integration of targeted exploration, online sample filtering, and experience replay, the framework not only addresses the limitations of GRPO in handling complex reasoning tasks but also establishes a more principled and controllable training paradigm. Moreover, by introducing a categorization of training samples based on their difficulty and informativeness, EFRame offers a new analysis of the roles of different samples within the RLVR framework. Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 36.5%, and it achieves consistent gains across a wide range of multimodal and mathematical reasoning benchmarks, demonstrating strong generalizability and robustness.
@@ 4.3 Online filter
Regular Filter: In the first-round rollout, when the prompt is of regular difficulty—we adopt a conservative filtering strategy that retains all regular samples $O_{R}$ to balance efficiency and performance.
@@ 4.4 Replay buffer
If the actor policy $\pi_{\theta}$ is a tabular softmax policy updated via natural policy gradient (Kakade, 2001) with step size $\eta$. Then the change in policy entropy under state between two steps approximately satisfies:
@@ 4.5 Discussion
Additional rollout: By prompting LLM to generate multiple responses to challenging queries ($x\in\bar{\mathcal{D}}$) under a high-temperature setting, the resulting positive samples are obtained with extremely low probability $\log\pi_{\theta}(y\mid x)$ and high advantage $\hat{A}(x,y)$. According to Theorem 4.5, this high-advantage, low-probability scenario rapidly increases the policy entropy $H$, thereby preventing the collapse of the entropy and maintaining effective exploration. At the same time, an excessive presence of low-quality samples can cause a nosedive of entropy, which is difficult to control. To address this, we filter out such samples and instead adopt a more controllable strategy to regulate entropy stability.
@@
Applying any techniques in isolation may result in unhealthy dynamics of policy entropy. Subsequent ablation studies further corroborate these findings.
@@ 5.1 Experimental setup
Implementation details. Our work is based on EasyR1 and VeRL framework Yaowei et al. (2025); Sheng et al. (2025). We initialize all policy models with Qwen2.5-VL-7B-Instruct and conduct experiments on 8 A800 GPUs (40GB). Unless otherwise specified, we follow the default EasyR1 settings: maximum response length of 2048, global batch size 128, rollout batch size 512, rollout $G_{1}$=5, temperature 1.0, and learning rate $10^{-6}$. GRPO and our method adopt $\epsilon_{low}=\epsilon_{high}=0.2$ and a binary reward to avoid relative advantage bias (see detail in Appendix A), while DAPO includes $\epsilon_{high}=0.3$ and an additional length-aware linear penalty (up to 0.5) for responses exceeding 1536 tokens. For our method, we perform additional rollouts with $G_{2}=20$, temperature = 1.2; responses with absolute advantage larger than 1 are stored in a replay buffer of size 5120, with experience replay triggered every 5 steps.
@@ 6 Ablation Studies
To assess the contribution of each component in our framework, we conduct an ablation study to examine whether all three modules—additional rollout, online filtering, and replay buffer—are all essential for achieving the observed improvements in stability and overall performance.
@@ 7 Conclusion and Discussion
In this work, we investigate the impact of sample quality on RL for large language models, with a particular focus on addressing challenging prompts. Our findings show that low-quality incorrect samples—those with low absolute advantage—tend to trap the model in local optima and offer little or even negative contribution to training. In contrast, high-quality samples with large advantages play a critical role in guiding the model through a meaningful transition from entropy-driven exploration to entropy-reducing convergence. This process is essential for eliciting deeper reasoning capabilities within the model. To this end, we propose an Exploration-Filter-Replay framework that emphasizes the discovery, selection, and sustained utilization of high-quality samples. Our method introduces additional rollouts to enhance reward density, employs online filtering to suppress noisy samples, and integrates experience replay to amplify the impact of high-quality samples across training iterations. This strategy not only helps transform isolated “aha moments” into a sustained “aha state,” but also shifts reinforcement learning from a “reinforcement lottery” toward a more structured and reliable process of reinforcement learning.
@@ Appendix A Relative Advantage Bias of GRPO
Appendix B Key Hyperparameters
The quantity of high-quality samples varies significantly with dataset difficulty, which makes our algorithm sensitive to three key hyperparameters: clip higher, replay buffer size, and replay frequency. This is because the gradients from high-quality samples must be just perfect—between underfitting, proper fitting, and overfitting—to be effective. Notably, on relatively simple datasets such as Math, where high-quality samples are inherently scarce, the improvements of our framework over the original GRPO are less pronounced.

### v3

@@ Abstract
Recent advances in reinforcement learning (RL) have significantly enhanced the reasoning capabilities of large language models (LLMs). Group Relative Policy Optimization (GRPO), an efficient variant of PPO that lowers RL’s computational cost, still faces limited exploration, low sample efficiency and instability, constraining its performance on complex reasoning tasks. To address these limitations, we introduce EFRame, an Exploration-Filter-Replay framework that systematically augments GRPO along three critical dimensions. EFRame performs additional rollouts to explore high-quality trajectories, applies online filtering to eliminate low-quality samples that introduce noise and variance, and leverages experience replay to repeatedly exploit rare but informative samples. EFRame establishes a complete and stable learning cycle, guiding the model through a structured transition from exploration to convergence. Our experiments across a variety of reasoning benchmarks demonstrate that EFRame not only improves the robustness and efficiency of training, but also enables access to deeper reasoning capabilities that remain unattainable under vanilla GRPO. Furthermore, EFRame not only enables fine-grained categorization of training samples for deeper insight into their contributions, but also introduces an efficient and precise mechanism for entropy control, which is critical for balancing exploration and convergence in RL training. Our code is available at https://github.com/597358816/EFRame.
@@ 1 Introduction
The contributions of EFRame are threefold:
  1. 1.
    Enhanced High-Difficulty Reasoning: EFRame significantly improves the performance of GRPO on challenging reasoning tasks. By integrating exploration, filtering, and replay into a unified framework, it enables more complete and deeper reasoning processes in large language models.
  2. 2.
    Sample-Type-Aware Training: EFRame introduces a principled categorization of training samples to better understand their distinct roles in learning. It demonstrates that high-advantage, low-probability positive samples are crucial for triggering exploration and unlocking deeper knowledge, whereas low-advantage negative samples tend to drive premature convergence to suboptimal solutions and should be suppressed.
  3. 3.
    Precise and Efficient Entropy Control: EFRame offers a practical and effective approach to entropy control by jointly tuning the additional rollout temperature and the number of replayed samples. This mechanism provides more flexible and stable entropy regulation than traditional reward-based entropy regularization, supporting both efficient exploration and stable convergence.
EFRame provides a systematic solution to the inherent trade-offs among training stability, efficiency, and exploration capacity in reinforcement learning. By enabling the model to effectively identify and leverage deeper latent knowledge, EFRame enhances the expressiveness and learning potential of large language models. Through the coordinated integration of targeted exploration, online sample filtering, and experience replay, the framework not only addresses the limitations of GRPO in handling complex reasoning tasks but also establishes a more principled and controllable training paradigm. Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 36.5%, and it achieves consistent gains across a wide range of multimodal and mathematical reasoning benchmarks, demonstrating strong generalizability and robustness.
@@ 4.3 Online filter
Regular Filter: In the first-round rollout, when the prompt is of regular difficulty, we adopt a conservative filtering strategy that retains all regular samples $O_{R}$ to balance efficiency and performance.
@@ 4.4 Replay buffer
If the actor policy $\pi_{\theta}$ is a tabular softmax policy updated via natural policy gradient (Kakade, 2001) with step size $\eta$. Then the change in policy entropy between two steps approximately satisfies:
@@ 4.5 Discussion
Additional rollout: By prompting LLM to generate multiple responses to challenging queries ($x\in\bar{\mathcal{D}}$) under a high-temperature setting, the resulting positive samples are obtained with extremely low probability $\log\pi_{\theta}(y\mid x)$ and high advantage $\hat{A}(x,y)$. According to Theorem 4.5, this high-advantage, low-probability scenario rapidly increases the policy entropy $H$, thereby preventing the collapse of the entropy and maintaining effective exploration. At the same time, an excessive presence of low-quality samples can cause a nosedive in entropy, which is difficult to control. To address this, we filter out such samples and instead adopt a more controllable strategy to regulate entropy stability.
@@ 5.1 Experimental setup
Implementation details. Our work is based on EasyR1 and VeRL frameworks Yaowei et al. (2025); Sheng et al. (2025). We initialize all policy models with Qwen2.5-VL-7B-Instruct and conduct experiments on 8 A800 GPUs (40GB). Unless otherwise specified, we follow the default EasyR1 settings: maximum response length of 2048, global batch size 128, rollout batch size 512, rollout $G_{1}$=5, temperature 1.0, and learning rate $10^{-6}$. GRPO and our method adopt $\epsilon_{low}=\epsilon_{high}=0.2$ and a binary reward to avoid relative advantage bias (see detail in Appendix A), while DAPO includes $\epsilon_{high}=0.3$ and an additional length-aware linear penalty (up to 0.5) for responses exceeding 1536 tokens. For our method, we perform additional rollouts with $G_{2}=20$, temperature = 1.2; responses with absolute advantage larger than 1 are stored in a replay buffer of size 5120, with experience replay triggered every 5 steps.
@@ 5.2 Main results
5.3 Entropy Control
To achieve a well-balanced exploration process, we incorporate entropy control into EFRame by adjusting two key hyperparameters: the additional rollout temperature $t_{a}$ and the replay buffer size $R_{s}$. Specifically, increasing the temperature in the additional rollout phase encourages broader exploration, enabling the model to discover low-probability but potentially high-reward trajectories. At the same time, we increase the $R_{s}$ to promote convergence and prevent excessive entropy growth. Together, these mechanisms allow for a controlled yet sufficiently rich exploration process—ensuring that the model explores meaningfully without destabilizing training due to overly random behavior.
((a))
((b))
((c))
((d))
((e))
((f))
Accuracy $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
$R_{s}=5120$ 53.74 52.41 55.24
$R_{s}=2560$ 52.75 54.41 52.75
$R_{s}=0$ 52.41 51.08 50.25
Entropy $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
$R_{s}=5120$ 0.089 0.305 0.726
$R_{s}=2560$ 0.091 0.312 2.366
$R_{s}=0$ 0.092 2.437 4.288
Figure 5: We report the peak accuracy within the first 200 training steps and the average entropy over the first 150 steps. Experiments on the GEO3K dataset using Qwen2.5-VL-7B-Instruct under different settings of additional rollout temperature $t_{a}$ and replay buffer size $R_{s}$ demonstrate that increasing $t_{a}$ enhances exploration, reflected in higher entropy, while enlarging $R_{s}$ promotes convergence, leading to lower entropy. Optimal performance is achieved when exploration and convergence are well-balanced.
We observe in Fig. 5 that increasing the additional rollout temperature encourages broader exploration, leading to a rise in policy entropy. In contrast, enlarging the replay buffer size, which increases the number of high-quality samples reused during training, promotes convergence and results in decreased entropy. These two hyperparameters exhibit opposing effects on entropy dynamics, providing a practical mechanism for entropy control within EFRame. We find that model performance peaks when the entropy stabilizes around a moderate level. Fig. 5 reports the accuracy under different $t_{a}$ and $R_{s}$, highlighting the impact of entropy control of EFRame on learning efficacy. Our experiments demonstrate that entropy control can be achieved with high precision by combining high-temperature positive sampling and experience replay, which are more effective than using entropy-based reward functions.
@@ 6 Ablation Studies
To assess the contribution of each component in our framework, we conduct an ablation study to examine whether all three modules—additional rollout, online filtering, and replay buffer—are essential for achieving the observed improvements in stability and overall performance.
@@ 7 Conclusion and Discussion
In this work, we investigate the impact of sample quality on RL for large language models, with a particular focus on addressing challenging prompts. Our findings show that low-quality incorrect samples—those with low absolute advantage—tend to trap the model in local optima and offer little or even negative contribution to training. In contrast, high-quality samples with large advantages play a critical role in guiding the model through a meaningful transition from entropy-driven exploration to entropy-reducing convergence. This process is essential for eliciting deeper reasoning capabilities within the model. To this end, we propose an Exploration-Filter-Replay framework that emphasizes the discovery, selection, and sustained utilization of high-quality samples. Our method introduces additional rollouts to enhance reward density, employs online filtering to suppress noisy samples, and integrates experience replay to amplify the impact of high-quality samples across training iterations. This strategy not only helps transform isolated “aha moments” into a sustained “aha state” but also shifts reinforcement learning from a “reinforcement lottery” toward a more structured and reliable process of reinforcement learning.